find all the zeros for f(x)=x to the fourth power - 6 times x to the third power + 9 times x to the second power + 6 times x - 10 if one zero is 3-i
Answer by stanbon(75887)

find all the zeros for f(x)=x to the fourth power - 6 times x to the third power + 9 times x to the second power + 6 times x - 10 if one zero is 3-i
-----------------
f(x) = x^4 - 6x^3 + 9x^2 + 6x - 10
---------------
If 3-i is a zero so is 3+i
Therefore (x-(3-i))(x-(3+i)) = ((x-3)+i)((x-3)+i)
= (x-3)^2+1 = x^2-6x+10 is a factor of f(x)
--------------
Using long division find the other factor(s):
The quotient is x^2-1, so the other zeroes are x=1 and x=-1
-------------------------------------
All the zeroes: 3-i,3+i,1,-1
==================
Cheers,
Stan H.